OpenStudy (anonymous):

Which of the following lines is not parallel to the graph of y = 9x + 1? 9y – x = 2 y – 9x = 3 9x – y = 2 27x – 3y = 2

OpenStudy (anonymous):

first put them in the same format y = mx + b like y = 9x + 1 m is the slope of the graph, in this case slope = 9 parallel lines have same slope

OpenStudy (anonymous):

y=x/9+2/9 y=9x+3 y=9x-2 y=9x-2/3

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

@nphuongsun93

OpenStudy (anonymous):

out of those 4 lines, which one doesn't have m = 9?

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

yup there u go

OpenStudy (anonymous):

thanks!